OpenStudy (anonymous):

Frank can complete a job in 5 hrs. Joe can do it in 8 hrs. After Frank has worked for 2 hrs. , Joe was asked to help Frank complete the job. In how many hours after Joe has started will they complete the job?

OpenStudy (amistre64):



OpenStudy (radar):

Well Frank has already got the job 2/5 done. That means only 3/5 of the job is left to do. Now joe and Frank together can do the complete job in (5)(8)/13=40/13 hours. But they only have 3/5 of the job left so it will take them (3/5)(40/13) or 24/13 hours. You could express that as a mixed number of 1 11/13 hour to complete or the entire job would of taken 2 + 1 11/13= 3 11/13. However the question asks How many hours after Joe starts helping to complete and that answer is 1 11/13 hrs.

OpenStudy (amistre64):